OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (8, 3 divided by 2 pi).

OpenStudy (tkhunny):

Have you considered the sine and cosine of the given angle?

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} \left(8,\frac{3}{2}\pi\right)\\ \ \uparrow\quad \uparrow \\ \ r\quad \theta \end{array}\qquad \begin{array}{llll} x=rcos(\theta)\\ y=rsin(\theta) \end{array}\qquad (x,y)\)

OpenStudy (anonymous):

so would it be (8,0) @jdoe0001

OpenStudy (tkhunny):

"3 divided by 2 pi" Is that \(\dfrac{3}{2\pi}\)? Sure sounds like it. Careful with notation. No, it's not (8,0). Think about \(3\pi/2\). Is that the positive y axis or negative y axis?